OpenStudy (anonymous):

Solve the equation 15^-x = 24. Round your answer to the nearest hundredth Solve the equation 15^-x = 24. Round your answer to the nearest hundredth @Mathematics

OpenStudy (anonymous):

\[-x=\frac{\ln(24)}{\ln(15)}\] \[x=-\frac{\ln(24)}{\ln(15)}\] then a calculator